OpenStudy (anonymous):

Dave can run the 440 yard dash in 55 seconds, and Jack can run it in 88 seconds. How great a handicap must Dave give Jack for the boys to finish the race at the same time?

OpenStudy (turingtest):

lol let me fix that

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

speed of dave=440/55=8 yard per sec speed of jack=440/88=5 yard per sec as dave must slow down to finish the race together as dave have to give a hand so must slow down

OpenStudy (anonymous):

so they must run with same speed= 5yard per sec so dave should slow down to 5 yard per sec so must slowed by 3 yard per sec